Fermi National Laboratory
 

Accelerator Update for Jun 7, 2006 - June 8, 2006

 

Wednesday June 7

The day shift began with Operations monitoring the Recycler stash and the Antiproton Source (Pbar) stack, with the Booster, Main Injector (MI), Recycler, and Switchyard (SY) conducting studies, and with NuMI, MiniBooNE, and the Tevatron (TeV) in access.

Operations prepared the Booster for an access at 10:18 AM.

Water technicians reported at 10:35 AM that they investigated the 400MeV Debuncher water skid and found nothing wrong. Increasing the head pressure must have fixed the problem.

At 10:48 AM, two MI RF stations (MIRF8 & 11) tripped off for the first of four times in forty-five minutes. They reset each time.

Operations reported the machine status at 11:16 AM:

  • The Recycle was stashing
  • TeV separators were conditioning
  • Transfer Hall access continued
  • EE-support investigated a LCW leak on the NuMI Horn power supply
  • Pbar Transport would be in access until noon
  • Booster in controlled access
  • F-sector controlled access on hold
  • D-Zero (D0) controlled access continues.
  • The Collider Detector at Fermilab (CDF) in access until 14:00 hrs
  • Meson M01 through M05 in access

Operations reported at 11:23 AM that the Recycler had been stashing antiprotons, but then the Recycler sequencer failed. Operators contacted two experts who were successful in fixing the problem.

Operations turned off F sector at 11:35 AM for an access.

The Recycler stash ended at 11:37 AM.

Water techs completed the NuMI Horn work at 11:41 AM.

Booster came out of access at 12:20 PM.

Operations reported at 12:54 PM that the Neutron Therapy Facility (NTF) completed patient treatment.

Booster experts established beam to the dump at 1:13 PM.

Techs came out of F sector at 2:15 PM.

Pbar came out of access at 2:40 PM.

Operations established beam to MiniBooNE at 2:59 PM.

At 3:03 PM, Operations reported that an Accumulator power supply (A:QT) had no output. When an expert investigated he discovered that it had been left breakered off after some work had been finished.

Operations established beam to NuMI at 3:29 PM.

CDF was out of access by 4:08 PM.

At 7:19 PM, A:QT tripped off on a ground fault indication. Operations called in an expert.

At 7:21 PM, the 400 MeV Debuncher cavity water flow tripped off for the first of three times in about thirty minutes.

Experts put the TeV into a dry squeeze at 9 PM.

At 9:41 PM, A:QT tripped off on a ground fault indication. It tripped off again forty-five minutes later.

D0 came out of access at 10:59 PM.

Operations put the TeV into a dry squeeze at 11:59 PM.

Thursday June 8

The midnight shift began with Operations monitoring the stash, with an expert investigating Pbar A:QT problems, with the TeV trying to come on, with MI, RR, and Switchyard (SY) conducting studies, with a Booster expert repairing a kicker, and with MiniBooNE and NuMI taking beam.

TeV experts began a study period at 12:22 AM.

A Booster expert repaired a sparking kicker (MKS11) at 12:30 AM.

At 1:07 AM, Operations reported that the 400 MeV Debuncher cavity water flow had tripped off six times in the last hour.

At 1:21 AM, the expert working on A:QT reported that he would continue work on it in the morning. He said he’d talk with the engineers in the morning. Pbar won’t be able to stack until this device gets repaired.

Between 6 and 7 AM, NuMI had many trips due a power supply’s (E:QD103) current drifting and causing the beam permit to drop out and due to BPM problems.

Experts resumed working on A:QT and at 7:35 AM had it up and running.

The Plans for Thursday

The plan is today is to get Pbar stacking and to send antiprotons to the TeV and then establish a small store tomorrow. The Run Coordinator would like the TeV to resume normal collider operations by next week.

Machine Reports

Linac

The machine has some LCW problems.

Booster

Experts replaced a PA for BRF1. BRF19 has a LCW leak. Experts say that they can run this station without cooling, but they will continue their investigations.

Main Injector

Experts accelerated antiprotons to 150 GeV. They have to work on the NuMI only cycle mode.

Antiproton Source

Experts reported that they had pulled out their prototype tank.

TeV

Experts report that both of their Tevatron Electron Lenses are working. Experts want to establish a 16x8 study-store on Friday. (16x8, the first number, 16, is for bunches of proton and the second number, 8, is for bunches of antiprotons.)

Recycler

Experts have cooled their beam and made antiproton transfers to MI.

Switchyard 120

The A0 quadrupole magnet was installed.

MiniBooNE

The experiment is taking data.

NuMI

Experts reported that they’re running okay, but at a low intensity.

Collider Detector at Fermilab

Experts found and fixed their noise and trigger problems. They took cosmic rays overnight.

D-Zero

Experts found their noise problem and made repairs, but there may be more problems.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/08/2006    email Fermilab